OpenStudy (anonymous):

What is the standard form of the circle given below? x^2+2x+y^2=55+10y

OpenStudy (anonymous):

A. (x-1)^2+(y-5^2)=81 B. (x+1)^2+(y-5)^2=9

OpenStudy (anonymous):

\[x^2+2x+y^2=55+10y\]\[x^2+2x+y^2-10y=55\]\[x^2+2x+1+y^2-10y+25=55+1+25\]\[(x+1)^2+(y-5)^2=81\]

OpenStudy (anonymous):

C. (x+1)^2+(y-5)^2=81

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

there is the answer, gave it to you.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

anytime